From mar@roe.ac.uk Mon Sep 2 08:48:54 2013 Date: Sun, 1 Sep 2013 16:40:14 +0100 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Alastair Edge , Andrew Lawrence , Bob Mann , Dante Minniti , James Dunlop , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Maria-Rosa Cioni , Matt Jarvis , Mike Irwin , Nigel Metcalfe , Norman Gray , Phil Lucas , Richard McMahon at IOA , Simon Dye , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives progress meeting minutes, 23 August 2013 Present: ETWS, MAR, NJC, RSC, RGM, MSH Apologies: NCH, STV, KTN, AL, JPE, RPB, ACD DoNM: 10am Friday 6th September 2013 in the Villas meeting room NCH in the chair (room is booked). Actions discharged this time: ---------------------------- ACTION: NJC to investigate the utility of the APOGEE Survey. It was decided not to host SDSS APOGEE at this time. Actions partly discharged but continuing: ---------------------------------------- ACTION: MAR to have a go at releasing other VISTA PI DBs (e.g. for JPE) Continues: JPE's programme release had been running but got hit by the problems with NAS box disks going offline. ACTION RPB and NJC: RPB to check the logs, identify the error and forward to NJC, who would investigate and fix it. Continues, this relates to a VISTA non-survey release that had fallen over. Actions carried forward from 9/8/2013 meeting: ---------------------------------------------- ACTION: All to think about items that could be included in the WFAU poster. ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this one still has low priority. Specific points and new actions: ------------------------------- Project management: The agenda for next Tuesday's WFAU Quarterly is on the Twiki. WFCAM, VISTA and VST updates: The University of Hawaii would take over the running of UKIRT at the end of the year. Comments and issues arising from CASU minutes: No new minutes this time. Networking: Working, though no new transfers due to the offline disks. WSA/VSA/OSA/GES Operations: Somewhat limited by the NAS box failure. ETWS reported: -- Created browser pages for the GES releases. -- Checking of VSA v1.3 catalogue ingests is ongoing (on hold due to the hardware problems). Hardware and Systems: MSH reported that ramses14 would soon be ready for VVV curation. However he noted that using RAID10 would only provide 39Tb of disk space. It was thought that the current VVV load server was already using that amount of storage but it was not clear if some of that could be recovered by eg shrinking the logs and any bloated filegroups. Using RAID5 would provide more but slower storage. ACTION: MSH/NJC to look and current and future VVV load server storage requirements so that RAMSES14 could be set-up optimally. MSH reported that the problem with the NAS box (Pami) was still being investigated but that the data were still on the disks and would be made available to the servers via the USB disk Linux installation. Eclipse were being consulted and when more was known a decision would be made about how to proceed. There was room on the new NAS box to allow a temporary copy of Pami's disks to be made if needed. Userkaf (MonetDB) was back up and running. Survey Data Release: The contents of the GESiDR1 release had been revised a couple of times on receipt of new/fixed data. Notes had been made on the release history. A revised VVV ESO release was being prepared. There was another discussion on the differences between WFAU and ESO VISTA releases. It is desirable for WFAU's world accessible VISTA releases to at least match or exceed the ESO releases in terms of content. ACTION: MAR/NJC to approach VISTA PIs about their public releases. Software: NJC reported further work on the matched aperture photometry pipeline. It's deployment was one of the issues discussed with Alastair Edge during his recent visit. ETWS reported: -- Installed Python 2.7.5 and recompiled 3rd party software relying on python. -- Working on the implementation of VISTA non-surveys into the schema parser. RSC reported: CU6 is now scaling nicely again for the VVV, with an overall 20x improvement in performance compared with our initial run, post the monthly-split of the detection tables. The code changes, having passed regression tests, have been merged back into trunk ready for DR2 operations to continue. It will take about a month to complete. The GES line list ingest code has been parallelised so that it completes in under an hour on our newer curation servers rather than taking more than a day. MAR mentioned work on additional functionality for the GES UI (getObject). Non-survey Data Release: Nothing to report. Astrogrid deployment, VO & Data Analysis services: All services running. Miscellaneous: Nothing of note. -- Scanned by iCritical.